My numerator is (4b-1)(4b+1) and the denominator is (2b-1)(b+6) how do I factor out my denominator to find the domain making sure I exclude the values equal to zero
Answer by Alan3354(69443)

My numerator is (4b-1)(4b+1) and the denominator is (2b-1)(b+6) how do I factor out my denominator to find the domain making sure I exclude the values equal to zero
-----------
Everything is factored.
---
For the DEN:
(2b-1)(b+6) = 0
b-6 = 0
--> b = -6 is excluded
===============
2b-1 = 0
b = 1/2 is excluded